# Parameter estimation in fluid flow models from undersampled frequency space data

### v1

@@ Abstract
4D Flow MRI is the state of the art technique for measuring blood flow, and it provides valuable information for inverse problems in the cardiovascular system. However, 4D Flow MRI has a very long acquisition time, straining healthcare resources and inconveniencing patients. Due to this, usually only a part of the frequency space is acquired, where then further assumptions need to be made in order to obtain an image.
Inverse problems from 4D Flow MRI data have the potential to compute clinically relevant quantities without the need for invasive procedures, and/or expanding the set of biomarkers for a more accurate diagnosis. However, reconstructing MRI measurements with Compressed Sensing techniques introduces artifacts and inaccuracies, which can compromise the results of the inverse problems. Additionally, there is a high number of different sampling patterns available, and it is often unclear which of them is preferable.
Here, we present a parameter estimation problem directly using highly undersampled frequency space measurements. This problem is numerically solved by a Reduced-Order Unscented Kalman Filter (ROUKF). We show that this results in more accurate parameter estimation for boundary conditions in a synthetic aortic blood flow than using measurements reconstructed with Compressed Sensing.
We also compare different sampling patterns, demonstrating how the quality of the parameter estimation depends on the choice of the sampling pattern. The results show a considerably higher accuracy than an inverse problem using velocity measurements reconstructed via compressed sensing. Finally, we confirm these findings on real MRI data from a mechanical phantom.
@@ 1 Introduction
In cardiovascular modeling in general and blood flows in particular, the personalization of spatially distributed (i.e. 3D) models is a key step in performing predictive patient-specific simulations [1]. This requires the estimation of relevant parameters from clinical data, which can be used in diagnostics or in the creation of patient-specific predictive simulations such as 3D models of the hemodynamics in the vascular system.
A common technique for acquiring measurements of blood flows is phase-contrast Magnetic resonance Imaging (MRI), or PC-MRI [2], as it is non-invasive and non-ionizing. However, acquiring 4D flow data with MRI requires long acquisition times, which strains clinical resources and inconveniences patients. This applies especially if a high spatial or temporal resolution in the data is required.
MRI applies sequences of magnetic fields and measures the precession of the quantum spin of hydrogen atoms in a strong magnetic field. The resulting measurement equates to the Fourier transformation of the complex magnetization, the phase of which contains information about the velocity. As such, MRI scans in frequency space (also called k-space). One of the features of MRI is that it is possible to select which frequencies of the k-space to measure. Therefore, to reduce the long acquisition times, often only a small part of the k-space is acquired [3].
This partial acquisition introduces artifacts in the reconstructed image that reduce both the image quality. As a result, various Compressed Sensing (CS) techniques have been developed, which aim to reconstruct velocities from highly undersampled data while minimizing artifacts [4]. CS has allowed for more important reductions in the sampling, in the sense that the accuracy of the reconstructed velocity images is higher than with the previously developed sampling and reconstruction strategies [5]. Nonetheless, as we will show later in this article, using CS-reconstructed data for inverse problems compromises the accuracy of the estimated flow and parameters considerably. Secondly, the potential of directly using the k-space measurements in the inverse problem has remain unexplored, while having the potential to avoid the introduction of reconstruction artifacts.
In this paper, we introduce a new technique for estimating blood flow parameters directly from the undersampled k-space MRI data by using a objective function designed for complex MRI measurements. This both avoids artifacts and improves efficiency by skipping the reconstruction/compressed sensing step.
@@ 2.1.1 Velocity encoding
Let us denote by $u(\bm{x},t)$ the component of the velocity $\bm{u}(\bm{x},t)$ in the direction $\bm{d}$ (fixed in space and in time), and $\bm{x}\in\mathbb{R}^{3}$, $t\in\mathbb{R}$ represent the spatial and temporal location within the image space, respectively. The MR images usually analyzed in the clinical context are actually complex valued, namely the so-called magnetization
$m(\bm{x},t)=M(\bm{x},t)e^{i\phi(\bm{x},t)}\in\mathbb{C},$ (1)
where the magnitude $M>0$ usually displays the anatomy, and the phase $\phi(\bm{x},t)\in(-\pi,\pi]$ in blood flow imagining takes the form
$\phi(\bm{x},t)=\frac{\pi}{venc}u(\bm{x},t)+\phi_{back}(\bm{x},t)$ (2)
@@
Since both $u$ and $\phi_{back}$ are unknown on every voxel, at least two measurements are required. In the simplest case (which is the most commonly used in clinical practice), one measurement with no gradient is acquired to obtain $\phi_{back}$, and a second one with $venc\neq 0$, leading to two complex magnetic measurements, so that the velocity can be simply obtained by subtracting the phases, namely,
@@
In practice, the image $m(\bm{x},t)$ is discrete in both space and time, being divided into voxels and time instants. A voxel refers to a spatial unit across which the signal is assumed to be constant to provide a single value. The number of voxels is determined by the spatial resolution. The time instants are the times at which a measurement is made, which are determined by the temporal resolution. Therefore $m(\bm{x},t)$ can be represented as a matrix in $\mathbb{C}^{N_{x}\times N_{y}\times N_{z}\times N_{T}}$, with $N_{x},N_{y},N_{z}$ being the number of voxels in each spatial direction and $N_{T}$ the number of time instants. We will refer to the total number of voxels as $N=N_{x}\cdot N_{y}\cdot N_{z}$.
@@ 2.1.2 The raw signal and image reconstruction
As a matter of fact, the raw signal measured by the MRI scanner corresponds to the spatial Fourier transform of the magnetization, which can, at each measured “time instant”, be formulated as
$\bm{Y}(\bm{k})=\mathcal{F}\left[\bm{M}\odot\exp\left(i\left(\frac{\pi}{venc}\bm{u}_{meas}+\phi_{back}\right)\right)\right](\bm{k})+\epsilon(\bm{k})$ (4)
where $\odot$ denotes the Hadamard product. We now consider discrete quantities $\bm{M}\in\mathbb{R}^{N}$, $\bm{u}_{meas}\in\mathbb{R}^{N}$ and $\mathcal{F}:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is the three-dimensional discrete Fourier transform defined as
$\mathcal{F}[\bm{X}](\bm{k})=\sum_{n_{x}=0}^{N_{x}-1}\sum_{n_{y}=0}^{N_{y}-1}\sum_{n_{z}=0}^{N_{z}-1}\bm{X}_{n_{x}n_{y}n_{z}}\exp\left(-i2\pi\left(\frac{k_{1}n_{x}}{N_{x}}+\frac{k_{2}n_{y}}{N_{y}}+\frac{k_{3}n_{z}}{N_{z}}\right)\right)dx$ (5)
and $\bm{\epsilon}(\bm{k})\in\mathbb{C}^{N}$ is a complex Gaussian noise with a mean of zero. Hence the measurements are $\bm{Y}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$.
In practice, for a given temporal resolution, only a number of k-space lines can be acquired. Therefore, the MRI scanner acquires different lines in k-space data at each cardiac cycles. Therefore a “time instant” $n$ actually contains frequencies measured during different cardiac cycles, under the assumption that there are no significant differences between consecutive cardiac cycles (or if there are differences, for instance in the heart rate, such data is rejected). There is also freedom in choosing different k-space trajectories at different instants of the cardiac cycle, though that is rarely done.
The noise level depends on the setup chosen in the scans. For instance, the lower the spatial resolution (i.e., the larger the voxel size), the higher the signal-to-noise ration (because the signal increases, for the same noise level of the device). Also, the lower the $venc$ the higher the sensitivity of the magnetization phase to the velocity, and therefore one expects lower noise in the reconstructed velocities with phase-contrast. As the phase is limited to the range $[-\pi,\pi)$, the velocity is however limited to the range $[-venc,venc)$. If the maximum velocity exceeds the chosen $venc$, this results in phase aliasing artifacts. As a result, a good choice of $venc$ necessitates a compromise between noise level and presence of artifacts.
If the frequency space is fully sampled, i.e. the values of all k-space locations are known, the velocity can be reconstructed by applying the inverse Fourier transform to $\bm{Y}^{n}$ and equation (3). However, fully sampling the k-space is not practicable in cardiovascular MRI scans since 3D scans may take of the order of hours, depending on the spatiotemporal resolution chosen.
@@
Some common 2D patterns include regular Cartesian sampling, radial sampling, and pseudo-spiral sampling [6]. These can be extended into 3D by stacking 2D designs, leading to the stack-of-stars and stack-of-spirals designs [6]. The patterns can differ in whether lines are arranged in a horizontal and vertical or in a shear pattern, in the number of radial lines or spiral arms, the number of points per line, and whether the patterns are stacked with equidistant or variable density in the z-direction. Additionally, with stack-of-stars and stack-of-spirals, there is the option of rotating the pattern by a certain angle for each level in the stack.
Other methods recommend to use pseudo-random sampling methods, usually with a higher density in the center, to ensure that the artifacts are incoherent and resemble noise. Here as well there are many choices for which probability distribution to pick the points from.
Undersampled data cannot be reconstructed with an inverse Fourier transform, as it violates the Nyquist limit. Compressed sensing (CS) techniques are commonly used to reconstruct undersampled data in a way that limits the presence of artifacts. These techniques often exploit the sparsity of the data in a different domain, e.g the wavelet domain. Generally this can be formulated as solving the optimization problem
$\bm{\hat{m}=}\operatorname*{\operatorname{argmin}}_{\bm{m}}\{||\mathcal{F}_{U}(\bm{m})-\bm{Y}||^{2}_{2}+\lambda||\Phi\bm{m}||_{1}\}$ (6)
where $\mathcal{F}_{U}$ is the Fourier transform combined with the undersampling, $\bm{Y}$ is the measured k-space data, $\bm{\hat{m}}$ is the reconstructed image data, $\Phi$ is a sparsifying transform, and $\lambda$ is the regularization parameter.
The sparsifying transform transports the image into a domain where the desired image is sparse. For MRI images, several such domains exist, for example wavelet domains or total variation time. A good choice of $\Phi$ depends on the spatial and temporal resolution as well as the nature of the artifacts. This makes pseudo-random masks appealing, as they lead to incoherent noise artifacts. The regularization parameter $\lambda$ is usually determined heuristically or empirically and it is known to have an impact in the reconstructed images.
@@ 2.2.1 Parameters estimation in PDEs
To consider the underlying physics of the the application, we assume that they can be modelled as the solution $\bm{u}:\mathbb{R}^{3}\times[0,T]\to\mathbb{R}^{3}$ of a partial differential equation
$F\left(\bm{u},\frac{\partial\bm{u}}{\partial x_{1}},\frac{\partial\bm{u}}{\partial x_{2}},\frac{\partial\bm{u}}{\partial x_{3}},\frac{\partial\bm{u}}{\partial t},\bm{\theta}\right)=0$ (7)
in a domain $\Omega\subset\mathbb{R}^{3}$ with an initial condition $\bm{u}(\bm{x},0)=\bm{u}^{0}$ and a set of boundary conditions which is dependent on a set of parameters $\bm{\theta}$, which can describe boundary conditions or material parameters. Then we can define the forward operator $\mathcal{A}(\bm{\theta})$ which describes the solution of this PDE for a vector of model parameters $\bm{\theta}\in\mathbb{R}^{p}$, which includes material parameters, values of boundary conditions, etc. As such, the forward problem generates data according to a physical model with given model parameters.
The goal of the inverse problem overlying this PDE is to retrieve a (sub)set of the parameters $\bm{\theta}$ given measurements of $\bm{u}(\bm{x},t)$.
@@ 2.2.2 Measurements
For formulating the inverse problem, it can be seen that we have three different choices for the measurements:
@@
    the frequency space data $\bm{Y}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$, which is complex-valued, raw data which might be undersampled. The noise on this data is Gaussian and zero-mean.
@@
    the magnetization data $\bm{M}_{meas}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$, which relates to the frequency data as $\bm{M}^{n}_{meas}=\mathcal{F}^{-1}(\bm{Y}^{n})=\bm{M}_{true}+\mathcal{F}^{-1}(\bm{\epsilon})$. This is also complex-valued data with a Gaussian noise with zero mean, as the inverse Fourier transform of a Gaussian is Gaussian as well. However, if the k-space is undersampled, this measurement has to be reconstructed with CS. In this case, the noise on the velocity measurements cannot assumed to be Gaussian and independent, as there are correlations between the noise in different pixels. [7]
@@
    the velocity data $\bm{V}^{n}\in\mathbb{R}^{N}$ for $n=1,\cdots,N_{T}$. $\bm{V}^{n}$ is calculated from the angle of the magnetization, meaning that in the case of undersampling the magnetization has to be reconstructed first using compressed sensing as well. Additionally, the noise on this measurement cannot be approximated well by a Gaussian distribution[8].

### v2

# Parameter Estimation in Fluid Flow Models from Highly Undersampled Frequency Space Data

@@ Abstract
4D Flow MRI is the state-of-the-art technique for measuring blood flow and provides valuable information for inverse problems in the cardiovascular system. However, 4D Flow MRI requires very long acquisition times, straining healthcare resources and inconveniencing patients. To address this, usually only a part of the frequency space is acquired, which necessitates further assumptions to obtain an image.
Inverse problems based on 4D Flow MRI data have the potential to compute clinically relevant quantities without invasive procedures and to expand the set of biomarkers for more accurate diagnosis. However, reconstructing MRI measurements with compressed sensing techniques introduces artifacts and inaccuracies, which can compromise the results of inverse problems. Additionally, there are many different sampling patterns available, and it is often unclear which is preferable.
Here, we present a parameter estimation problem that directly uses highly undersampled frequency space measurements. This problem is numerically solved by a Reduced-Order Unscented Kalman Filter (ROUKF). We show that this approach results in more accurate parameter estimation for boundary conditions in a synthetic aortic blood flow model than using measurements reconstructed with compressed sensing.
We also compare different sampling patterns, demonstrating how the quality of parameter estimation depends on the choice of sampling pattern. The results show considerably higher accuracy than inverse problems using velocity measurements reconstructed via compressed sensing. Finally, we confirm these findings on real MRI data from a mechanical phantom.
Keywords: 4D MRI, inverse problems, Kalman filter, finite element method
@@ 1 Introduction
In cardiovascular modeling in general and blood flows in particular, the personalization of spatially distributed (i.e. 3D) models is a key step in performing predictive patient-specific simulations [nolte2022review]. This requires the estimation of relevant parameters from clinical data, which can be used in diagnostics or in the creation of patient-specific predictive simulations such as 3D models of the hemodynamics in the vascular system. A common technique for acquiring measurements of blood flows is phase-contrast Magnetic Resonance Imaging (MRI), or PC-MRI [markl2003jmri], as it is non-invasive and non-ionizing. However, acquiring 4D flow data with MRI requires long acquisition times, which strains clinical resources and inconveniences patients. This applies especially if a high spatial or temporal resolution in the data is required.
MRI applies sequences of magnetic fields and measures the precession of the quantum spin of hydrogen atoms in a strong magnetic field. The resulting measurement equates to the Fourier transformation of the complex magnetization, the phase of which contains information about the velocity. As such, MRI scans in frequency space (also called k-space). One of the features of MRI is that it is possible to select which frequencies of the k-space to measure. Therefore, to reduce the long acquisition times, often only a small part of the k-space is acquired [gottwald2020pseudo].
This partial acquisition introduces artifacts in the reconstructed image that reduce the image quality. As a result, various Compressed Sensing (CS) techniques have been developed, which aim to reconstruct velocities from highly undersampled data while minimizing artifacts. CS has allowed for more important reductions in the sampling, in the sense that the accuracy of the reconstructed velocity images is higher than with the previously developed sampling and reconstruction strategies [peper2020highly, cs2007, ye_compressed_2019].
Nonetheless, as we will show later in this article, using CS-reconstructed data for parameter estimation compromises the accuracy of the estimated flow and parameters considerably. Secondly, the potential of directly using the k-space measurements in the inverse problem to estimate parameters in blood flow has remained unexplored, while having the potential to avoid the introduction of reconstruction artifacts.
In this paper, we introduce a new technique for estimating blood flow parameters directly from the undersampled k-space MRI data by using an objective function designed for complex MRI measurements. This both avoids artifacts and improves efficiency by skipping the reconstruction/compressed sensing step. Preliminary results in this direction have been shown in [locke_parameter_2025-1] without being explored in-depth. This paper provides additional notes on the methodology as well as further results for different $V_{enc}$ values and the effects per parameter, as well as exploring the robustness of the method to using approximations of the magnitude and background phase. Furthermore, we are also including results on real MRI phantom data.
@@ 2.1.1 Velocity encoding
Let us denote by $u(\boldsymbol{x},t)$ the component of the velocity $\boldsymbol{u}(\boldsymbol{x},t)$ in the direction $\boldsymbol{d}$ (fixed in space and in time), and $\boldsymbol{x}\in\mathbb{R}^{3}$, $t\in\mathbb{R}$ represent the spatial and temporal location within the image space, respectively. The MR images usually analyzed in the clinical context are actually complex valued, namely the so-called magnetization
$m(\boldsymbol{x},t)=M(\boldsymbol{x},t)\exp{(i\phi(\boldsymbol{x},t))}\in\mathbb{C},$ (1)
where the magnitude $M>0$ usually displays the anatomy, and the phase $\phi(\boldsymbol{x},t)\in(-\pi,\pi]$ in blood flow imagining takes the form
$\phi(\boldsymbol{x},t)=\frac{\pi}{venc}u(\boldsymbol{x},t)+\phi_{back}(\boldsymbol{x},t)$ (2)
@@
Since both $u$ and $\phi_{back}$ are unknown at every voxel, at least two measurements are required. In the simplest case (which is the most commonly used in clinical practice), one measurement with no velocity encoding gradient is acquired to obtain $\phi_{back}$, and a second one with $venc\neq 0$, leading to two complex magnetic measurements, so that the velocity can be simply obtained by subtracting the phases, namely,
@@
In practice, the image $m(\boldsymbol{x},t)$ is discrete in both space and time, being divided into voxels and time instants. A voxel refers to a spatial unit across which the signal is assumed to be constant to provide a single value. The number of voxels is determined by the spatial resolution. The time instants are the times at which a measurement is made, which are determined by the temporal resolution. Therefore $m(\boldsymbol{x},t)$ can be represented as a matrix in $\mathbb{C}^{N_{x}\times N_{y}\times N_{z}\times N_{T}}$, with $N_{x},N_{y},N_{z}$ being the number of voxels in each spatial direction and $N_{T}$ the number of time instants. We will refer to the total number of voxels as $N=N_{x}\cdot N_{y}\cdot N_{z}$.
@@ 2.1.2 The raw signal and image reconstruction
The raw signal measured by the MRI scanner corresponds to the spatial Fourier transform of the magnetization, which can, at each measured “time instant”, be formulated as
$\boldsymbol{Y}(\boldsymbol{k})=\mathcal{F}\left[\boldsymbol{M}\odot\exp\left(i\left(\frac{\pi}{venc}\boldsymbol{u}_{meas}+\phi_{back}\right)\right)\right](\boldsymbol{k})+\epsilon(\boldsymbol{k})$ (4)
where $\odot$ denotes the Hadamard product. We now consider discrete quantities $\boldsymbol{M}\in\mathbb{R}^{N}$, $\boldsymbol{u}_{meas}\in\mathbb{R}^{N}$ and $\mathcal{F}:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is the three-dimensional discrete Fourier transform defined as
$\mathcal{F}[\boldsymbol{X}](\boldsymbol{k})=\sum_{n_{x}=0}^{N_{x}-1}\sum_{n_{y}=0}^{N_{y}-1}\sum_{n_{z}=0}^{N_{z}-1}\boldsymbol{X}_{n_{x}n_{y}n_{z}}\exp\left(-i2\pi\left(\frac{k_{1}n_{x}}{N_{x}}+\frac{k_{2}n_{y}}{N_{y}}+\frac{k_{3}n_{z}}{N_{z}}\right)\right)dx$ (5)
and $\boldsymbol{\epsilon}(\boldsymbol{k})\in\mathbb{C}^{N}$ is a complex Gaussian noise with a mean of zero[IRARRAZAVAL2019]. Hence the measurements are $\boldsymbol{Y}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$.
In practice, for a given temporal resolution, only a number of k-space lines can be acquired. Therefore, the MRI scanner acquires different lines in k-space during each cardiac cycle. Therefore a “time instant” $n$ actually contains frequencies measured during different cardiac cycles, under the assumption that there are no significant differences between consecutive cardiac cycles (or if there are differences, for instance in the heart rate, such data is rejected). There is also freedom in choosing different k-space trajectories at different instants of the cardiac cycle, though that is rarely done.
The noise level depends on the setup chosen in the scans. For instance, the lower the spatial resolution (i.e., the larger the voxel size), the higher the signal-to-noise ratio (because the signal increases, for the same noise level of the device). Also, the lower the vencvenc the higher the sensitivity of the magnetization phase to the velocity, and therefore one expects lower noise in the reconstructed velocities with phase-contrast. As the phase is limited to the range $\pi,\pi)$, the velocity is however limited to the range [venc,venc)[-venc,venc). If the maximum velocity exceeds the chosen vencvenc, this results in phase aliasing artifacts. As a result, a good choice of vencvenc necessitates a compromise between noise level and presence of artifacts.
Undersampling and reconstruction. If the frequency space is fully sampled, i.e. the values of all k-space locations are known, the velocity can be reconstructed by applying the inverse Fourier transform to $\boldsymbol{Y}^{n}$ and equation (3). However, fully sampling the k-space is not practicable in cardiovascular MRI scans since 3D scans may take of the order of hours, depending on the spatiotemporal resolution chosen.
@@
Some common 2D patterns include regular Cartesian sampling, radial sampling, and pseudo-spiral sampling [feng2022review]. These can be extended into 3D by stacking 2D designs, leading to the stack-of-stars and stack-of-spirals designs [feng2022review]. The patterns can differ in whether lines are arranged in a horizontal and vertical or in a shear pattern, in the number of radial lines or spiral arms, the number of points per line, and whether the patterns are stacked with equidistant or variable density in the z-direction. Additionally, with stack-of-stars and stack-of-spirals, there is the option of rotating the pattern by a certain angle for each level in the stack.
Another option is to use pseudo-random sampling methods, usually with a higher density in the center, to ensure that the artifacts are incoherent and resemble noise [peper2020highly]. Here as well there are many choices for which probability distribution to use.
Undersampled data cannot be reconstructed with an inverse Fourier transform, as it violates the Nyquist limit. Compressed sensing (CS) techniques are commonly used to reconstruct undersampled data in a way that limits the presence of artifacts. Generally this can be formulated as solving the optimization problem
[…]
where $\mathcal{F}_{U}$ is the Fourier transform combined with the undersampling, $\boldsymbol{Y}$ is the measured k-space data, $\boldsymbol{\hat{m}}$ is the reconstructed image data, $\Phi$ is a sparsifying transform, and $\lambda$ is the regularization parameter.
The sparsifying transform transports the image into a domain where the desired image is sparse. For MRI images, several such domains exist, for example wavelet domains or total variation in time. A good choice of $\Phi$ depends on the spatial and temporal resolution as well as the nature of the artifacts. This makes pseudo-random masks appealing, as they lead to incoherent noise artifacts. The regularization parameter $\lambda$ is usually determined heuristically or empirically and is known to have an impact on the quality of the reconstructed images.
@@ 2.2.1 Parameters estimation in PDEs
To consider the underlying physics of the the application, we assume that they can be modelled as the solution $\boldsymbol{u}:\mathbb{R}^{3}\times[0,T]\to\mathbb{R}^{3}$ of a partial differential equation
$F\left(\boldsymbol{u},\frac{\partial\boldsymbol{u}}{\partial x_{1}},\frac{\partial\boldsymbol{u}}{\partial x_{2}},\frac{\partial\boldsymbol{u}}{\partial x_{3}},\frac{\partial\boldsymbol{u}}{\partial t},\boldsymbol{\theta}\right)=0$ (7)
in a domain $\Omega\subset\mathbb{R}^{3}$ with an initial condition $\boldsymbol{u}(\boldsymbol{x},0)=\boldsymbol{u}^{0}$ and a set of boundary conditions which is dependent on a set of parameters $\boldsymbol{\theta}$, which can describe boundary conditions or material parameters. Then we can define the forward operator $\mathcal{A}(\boldsymbol{\theta})$ which describes the solution of this PDE for the vector of model parameters $\boldsymbol{\theta}\in\mathbb{R}^{p}$. As such, the forward problem generates data according to a physical model with given model parameters.
The goal of the inverse problem overlying this PDE is to estimate a (sub)set of the parameters $\boldsymbol{\theta}$ given measurements of $\boldsymbol{u}(\boldsymbol{x},t)$.
@@ 2.2.2 Measurements
For formulating the inverse problem to estimate the parameters $\boldsymbol{\theta}$, we have three different choices for the measurements:
@@
    the frequency space data $\boldsymbol{Y}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$, which is complex-valued, raw data which might be undersampled. The noise on this data is Gaussian, independently distributed, and zero-mean.
@@
    the magnetization data $\boldsymbol{M}_{meas}^{n}\in\mathbb{C}^{N}$ for $n=1,\cdots,N_{T}$, which relates to the frequency data as $\boldsymbol{M}^{n}_{meas}=\mathcal{F}^{-1}(\boldsymbol{Y}^{n})=\boldsymbol{M}_{true}+\mathcal{F}^{-1}(\boldsymbol{\epsilon})$. This is also complex-valued data with a Gaussian noise with zero mean, as the inverse Fourier transform of a Gaussian is Gaussian as well. However, if the k-space is undersampled, this measurement has to be reconstructed with CS. In this case, the noise on the velocity measurements cannot assumed to be Gaussian and independently distributed, as there are correlations between the noise in different pixels. [partin2022analysisreconstructionnoiseundersampled]
@@
    the velocity data $\boldsymbol{V}^{n}\in\mathbb{R}^{N}$ for $n=1,\cdots,N_{T}$. $\boldsymbol{V}^{n}$ is calculated from the angle of the magnetization, meaning that in the case of undersampling the magnetization has to be reconstructed first using compressed sensing as well. Additionally, the noise on this measurement cannot be approximated well by a Gaussian distribution[IRARRAZAVAL2019].